To borrow $1,300, you are offered an add on interest loan at 8.8 percent with 12 monthly payments. Compute the 12 equal payments.

Use the amount you borrowed and the monthly payments you computed to calculate the APR of the loan. Then, use that APR to compute the EAR of the loan.
Answer #1

Payment=(1300*1.088)/12=117.8666667
APR=RATE(12,-117.8666667,1300)*12=15.86%
EAR=(1+15.86%/12)^12-1=17.07%